Convert Log (a+b) 200 to the base 2

Convert Log(a+b)200 to the base 2

Grade:

2 Answers

Vijay Luxmi Askiitiansexpert
14 years ago

Hi Vivek ,

Convert  Log(a+b)200 to the base 2.

 

                                         Log(a+b)200   =  Log2200 / Log2(a+b)

The numerator can be calculated by converting the numerator to the base 10 or the base e.

                                 

                                      Log2200  =  ln200/ln2  =  7.6438562

                               The answer  Log2200 / Log2(a+b)  can also be written  7.6438562/Log2(a+b) .

Note that for this problem to be valid, the base (a + b) must be greater than zero.

pratik nayak
13 years ago

hey vivek

 

                                         Log(a+b)200   =  Log2200 / Log2(a+b)

the above expression comes from change of base theorem